OpenStudy (anonymous):

solve for x: 3|x-3|+2=14 A)No solution B)x=-1,x=8.3 C)x=0,x=7 D)x=-1,x=7

pooja195 (pooja195):

Add -2 to both sides. \[\huge~\rm~3(|x−3|)+2+−2=14+−2\] \[\huge~\rm~3(|x−3|)=12\] Then we need to divide (both sides) \[\huge~\rm~\frac{ 3(|x−3|) }{ 3 }=\frac{ 12 }{ 3}\] We get \[\huge~\rm~|x-3|=4\] Solve Absolute Value. \[\huge~\rm~|x−3|=4\] We know either :\[\huge~\rm~x−3=4\] \(\huge\color{red}{or}\) \[\huge~\rm~x−3=−4\]

pooja195 (pooja195):

From here just solve for x

OpenStudy (anonymous):

so i got x=7 and x=-1

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Thats correct! :) Good job! welcome to openstudy! :D

OpenStudy (anonymous):

Thank you so much for your help :D

pooja195 (pooja195):

No problem! :D